OpenStudy (anonymous):

A floor plan is given for the first floor of a new house. One inch represents 10 feet. If the walk-in closet is 2/5 in. by 2/3 in. on the floor plan, what are the actual dimensions? (Round to the nearest hundredth)

OpenStudy (anonymous):

if 1 inch is the same as 10 feet, then the ratio is: 1:10

OpenStudy (anonymous):

then seeing as the walk-in closet is 2/5 in. by 2/3 in. then use this ration to find the dimension, so: \[\frac{ 1 }{ 10 }=\frac{ { 2/5 } }{ x }\] solve for x, and do the same for the other length

jimthompson5910 (jim_thompson5910):

1 inch : 10 ft (2/5)*1 inch : (2/5)*10 ft ... multiply both parts by 2/5 2/5 inch : 20/5 ft 2/5 inch : 4 ft Do the same for the other dimension.

OpenStudy (anonymous):

Do i use cross multiplication

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 4 is the first answer

jimthompson5910 (jim_thompson5910):

2/5 inches on the map corresponds to 4 feet, yes

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

just need to find out what 2/3 inches corresponds to

OpenStudy (anonymous):

I need some help with the 2/3 part

jimthompson5910 (jim_thompson5910):

if you use the method I used, then you would start with this 1 inch : 10 ft and multiply both parts by 2/3

OpenStudy (anonymous):

well for the other one i turned 2/5 into .4 so it would be easier for me to multiply

jimthompson5910 (jim_thompson5910):

you can do that but keep in mind that 2/3 = 0.66667 so you kinda lose some info about the number 2/3

OpenStudy (anonymous):

for the second number would it be 6.67?

jimthompson5910 (jim_thompson5910):

yep, which is 20/3 if you chose to keep everything as fractions

jimthompson5910 (jim_thompson5910):

but 6.67 feet is probably easier to visualize

OpenStudy (anonymous):

TY!!!

jimthompson5910 (jim_thompson5910):

np